3. Express f(x) = x2 + 4x +3 in standard form and find its maximum or minumum value.
Answer #1

Griven f(2)= 27 4x+3 In Standard form x +41+3=0 According to Rolles theorem in order to find madison or minimum Finde .com) l(+48+3) = 0 2x + 4 = 0 2x=-4 at x=-2 it has minimum value substitute in I get minimerm volue. (2)+44(-2) +3 » 4- 873 to a so,